OpenStudy (anonymous):

What is the slope of x + 4y = 8

OpenStudy (anonymous):

y=mx+b 4y=-x+8 y= -x+8/4 the y int is m so y int = -1

OpenStudy (anonymous):

slope is \[-\frac{1}{4}\]

OpenStudy (anonymous):

\[x+4y=8\] \[4y=-x+8\] \[y=-\frac{1}{4}x+2\]